OpenStudy (anonymous):

everybody Doing partial fraction decomposition and need help finding the answer to this using equating coefficents 3x-8=A(x-1)+B(x-5)

OpenStudy (anonymous):

put \(x=1\) get \[-5=-4B\] pretty much instantly

OpenStudy (anonymous):

but i want to solve it the long way by equating coefficients and i don't know how

OpenStudy (anonymous):

that is the simple way to do it, equatiing like coefficients is the donkey way, but we can do that if you like

OpenStudy (anonymous):

yes do do :D

OpenStudy (anonymous):

\[3x-8=A(x-1)+B(x-5)\] multiply out on the right \[3x-8=Ax-A+Bx-5B\]

OpenStudy (anonymous):

combine like terms on the right \[3x-8=(A+B)x-A-5B\]

OpenStudy (anonymous):

that means \[A+B=3\\ -A-5B=-8\]

OpenStudy (anonymous):

satellite come back to my aueastion i am not sure

OpenStudy (anonymous):

then solve that system of equations add and get \[-4B=-5\] which (hate to say i told you so) i got in one second in my head in line one above

OpenStudy (anonymous):

pleaase

OpenStudy (anonymous):

thank you satellite :)

OpenStudy (anonymous):

yw